Two systems have thermodynamic probabilities of 3.0 * 10^27 and 1.8 10*28 respectively calculate the entropies of the individual systems as well as their in composite systems and verify the boltzsmen relation

1
Expert's answer
2022-04-17T17:06:51-0400

"S_1=\frac{s_2}{s_1+s_2}=0.86,"

"S_2=\frac{s_1}{s_1+s_2}=0.14,"

"S_1+S_2=1."